The following sample observations were randomly selected. (Round intermediate calculations and final answers to 2 decimal places.) 4 6 4 &Click here for the Excel Data File a. The regression equation is y- When X is 5 this gives y =
Answer #1

For the given set of data

a) The linear regression analysis is done by using calculation as

Calculation Summary

Sum of X = 19
Sum of Y = 24
Mean X = 3.8
Mean Y = 4.8
Sum of squares (SSX) = 8.8
Sum of products (SP) = -8.2

Regression Equation = ŷ = bX + a

b = SP/SSX = -8.2/8.8 = -0.93182

a = MY - bMX = 4.8 - (-0.93*3.8) = 8.34091

ŷ = -0.93182X + 8.34091

b) According to equation when equation for linear regression line is

ŷ = -0.93182X + 8.34091 hence at X= 5

Y =3.68